OpenStudy (anonymous):

Find the percent increase from 330 to 480

OpenStudy (anonymous):

45.5%

OpenStudy (anonymous):

\[\frac{330}{1}=\frac{480}{x}\] cross multiply \[330x=480\] divide by 330 \[\frac{480}{330}\approx 1.455\] multiply by 100 to get percentage \[1.455*100=145.5\]

OpenStudy (anonymous):

then minus 100 to get how much it increased

OpenStudy (anonymous):

{480-330/(330)}*100% =45.455%

OpenStudy (anonymous):

simple method is the above 1 i did